What would be the solution be to (13^1/20)^1/2
Answer by jmichau3(2)

not sure what you want to do with this, but you can simplify your exponents...
for example
(x^2)^3 = x^6
so for you, (13^1/20)^1/2 = (13^1/40) = 1.066224319
NOt sure if that is what you are looking for or not....hope it helps